# Text Indexing and Pattern Matching with Ephemeral Edits

### v1

@@ 5.2 Edit Operation
Since we have edit operations involving a single letter, it is readily verifiable that only Type 1, 2, 5, and 6 occurrences are relevant. Assume that we insert letter $c$ after $T[p]$. If $|P|=1$, a Type 6 occurrence is trivially reported in $\mathcal{O}(1)$ time. For the occurrences of $P$ in $T[0\mathinner{.\,.}p)$ (Type 1) we use RmQs and for the occurrences of $P$ in $T[p+1\mathinner{.\,.}|T|)$ (Type 2) we use RMQs, similar to Section 3. Namely, each occurrence is reported in $\mathcal{O}(1)$ time. For Type 5, we find the longest suffix of $P$ occurring as a prefix of $cT[p+1\mathinner{.\,.}n)$ as follows. When $p+1$ is given, we find the lex-rank $r=\textsf{iSA}[p+1]$ of $T[p+1\mathinner{.\,.}n)$ among all suffixes of $T$. We find the leaf $v_{p+1}$ representing $T[p+1\mathinner{.\,.}n)$ and follow the pointer to its deepest marked ancestor $u_{p+1}$. From there, using letter $c$, we find the longest suffix $cP[i\mathinner{.\,.}|P|)$ of $P$ occurring as a prefix of $cT[p+1\mathinner{.\,.}n)$ in $\mathcal{O}(1)$ time as follows. We follow the pointer of $u_{p+1}$ to the SMA. From the state corresponding to the suffix of $P$, represented by $u_{p+1}$, we follow the transition (if any) using letter $c$. This will bring us to the state that represents the reverse of $cP[i\mathinner{.\,.}|P|)=P[i-1]P[i\mathinner{.\,.}|P|)$. Recall that to find the longest prefix of $P$ that is a suffix of $T[0\mathinner{.\,.}p]$ we do the same work as for the deletion operation on the reverse of $P$ and $T$. After finding the longest prefix of $P$ that is a suffix of $T[0\mathinner{.\,.}p]$ and the longest suffix of $P$ occurring as a prefix of $cT[p+1\mathinner{.\,.}n)$, we trigger one prefix-suffix query. The total time to process the insertion is $\mathcal{O}(1)$ and for reporting is $\mathcal{O}(\textsf{Occ})$. The algorithm is correct by 2.
Recall that substitutions work the same way as insertions, but with a different offset. We have arrived at Theorem 3.
@@ 6 Concluding Remarks
It would be interesting to improve Theorem 1 by improving the preprocessing cost for the pattern or improving the query time for reporting. It would also be interesting to generalize Theorem 3 to support short substring insertions or substitutions. Another direction would be to consider other versions of text indexing or pattern matching (here we have considered their vanilla version) with ephemeral edits.

### v2

@@ 5.2 Edit Operation
For simplicity, let us start by edit operations involving a single letter; it is readily verifiable that only Type 1, 2, 5, and 6 occurrences are relevant. Assume that we insert letter $c$ after $T[p]$. If $|P|=1$, a Type 6 occurrence is trivially reported in $\mathcal{O}(1)$ time (otherwise, this type of an occurrence is not relevant). For the occurrences of $P$ in $T[0\mathinner{.\,.}p]$ (Type 1) we use RmQs and for the occurrences of $P$ in $T[p+1\mathinner{.\,.}|T|)$ (Type 2) we use RMQs, similar to Section 3. Namely, each occurrence is reported in $\mathcal{O}(1)$ time. For Type 5, we find the longest suffix of $P$ occurring as a prefix of $cT[p+1\mathinner{.\,.}|T|)$ as follows. When $p+1$ is given, we find the lex-rank $r=\textsf{iSA}[p+1]$ of $T[p+1\mathinner{.\,.}|T|)$ among all suffixes of $T$. We find the node $v_{p+1}$ representing $T[p+1\mathinner{.\,.}|T|)$ in $\textsf{ST}(T\#P)$ and follow the pointer to its deepest marked ancestor $u_{p+1}$. From there, using letter $c$, we find the longest suffix $cP[i\mathinner{.\,.}|P|)$ of $P$ occurring as a prefix of $cT[p+1\mathinner{.\,.}|T|)$ in $\mathcal{O}(1)$ time as follows. We follow the pointer of $u_{p+1}$ to the SMA. From the state corresponding to the suffix of $P$, represented by $u_{p+1}$, we follow the transition (if any) using letter $c$. This brings us to the state that represents the reverse of $cP[i\mathinner{.\,.}|P|)=P[i-1]P[i\mathinner{.\,.}|P|)$. Recall that to find the longest prefix of $P$ that is a suffix of $T[0\mathinner{.\,.}p]$ we do the same work as for the deletion operation on the reverse of $P$ and $T$. After finding the longest prefix of $P$ that is a suffix of $T[0\mathinner{.\,.}p]$ and the longest suffix of $P$ occurring as a prefix of $cT[p+1\mathinner{.\,.}n)$, we trigger one prefix-suffix query. The total time to process the insertion is $\mathcal{O}(1)$ and for reporting is $\mathcal{O}(\textsf{Occ})$. The algorithm is correct by 2.
@@
When handling a string $S$ of length $|S|=\mathcal{O}(1)$ (instead of a single letter $c$), the only difference for Type 5 is that from the SMA state corresponding to the suffix of $P$, we follow the transition path using the reverse of $S$. This brings us to the state that represents the reverse of $S\cdot P[i\mathinner{.\,.}|P|)$. This is correct because, for any strings $U$ and $X$, the $\textsf{SMA}(X)$ takes us to the longest prefix of $X$ that is a suffix of $U$. Thus, when $X$ is the reverse of $P$, and $U$ is the reverse of $S\cdot T[p+1\mathinner{.\,.}|T|)$, the SMA takes us to the longest suffix of $P$ that is a prefix of $S\cdot T[p+1\mathinner{.\,.}|T|)$. Type 3, Type 4, and Type 6 occurrences are handled exactly as in Section 3. Finally, recall that substitutions work the same way as insertions, but with a different offset. We have arrived at Theorem 3.
@@ 6 Concluding Remarks
It would be interesting to improve Theorem 1 by improving the preprocessing cost for the pattern or improving the query time for reporting. Another direction would be to consider other versions of text indexing or pattern matching with ephemeral edits (here we have considered their vanilla version).